# MuseumMaker: Continual Style Customization without Catastrophic Forgetting

### v1

@@ Abstract
Pre-trained large text-to-image (T2I) models with an appropriate text prompt has attracted growing interests in customized images generation field. However, catastrophic forgetting issue make it hard to continually synthesize new user-provided styles while retaining the satisfying results amongst learned styles. In this paper, we propose MuseumMaker, a method that enables the synthesis of images by following a set of customized styles in a never-end manner, and gradually accumulate these creative artistic works as a Museum. When facing with a new customization style, we develop a style distillation loss module to transfer the style of the whole dataset into generation of images. It can minimize the learning biases caused by content of images, and address the catastrophic overfitting issue induced by few-shot images. To deal with catastrophic forgetting amongst past learned styles, we devise a dual regularization for shared-LoRA module to optimize the direction of model update, which could regularize the diffusion model from both weight and feature aspects, respectively. Meanwhile, a unique token embedding corresponding to this new style is learned by a task-wise token learning module, which could preserve historical knowledge from past styles with the limitation of LoRA parameter quantity. As any new user-provided style come, our MuseumMaker can capture the nuances of the new styles while maintaining the details of learned styles. Experimental results on diverse style datasets validate the effectiveness of our proposed MuseumMaker method, showcasing its robustness and versatility across various scenarios.
@@ I Introduction
Although most recent models[8][9] focus on boosting their generative capabilities with natural languages, these models fail to control the exact texture and color palettes of generation images, when adding popular or customized styles (e.g., “impressionist”) to the input text prompt. A naive way is to fine-tune the T2I diffusion model using few images of the given “impressionist” style, whereas this manner cannot synthesize various images of different specific styles in a never-ending manner. For example, as shown in Fig. 1, when the user inputs a prompt such as “a cat wearing sunglasses in the style of impressionism” to generate an impressionist image of the cat, a naive way is to fine-tune the diffusion model with images in impressionist style. Subsequently, if the user tries to augment this diffusion model with a new style, e.g., “cubism”, a simple method is to fine-tune the diffusion model with both impressionistic and cubist images. As users continuously seek to incorporate new styles, the large computational requirements and ever-increasing training times swiftly become impractical. Moreover, there is a high probability for diffusion model to overfit to the new user-provided styles after continual fine-tuning operator, resulting in unsatisfactory artistic generation performance.
@@
Inspired by the aforementioned considerations, we present a continual style learning approach based on T2I diffusion model in this work. Our proposed method aims to enable the continuous learning of new styles for generation purposes without accessing past datasets, which is designed to reduce memory consumption and compute efficiently. To address the two aforementioned challenges, we develop a continual style customization for diffusion model i.e, MuseumMaker, which could continuously adapt upcoming new styles and purely learn stylistic features from user-provided data. To tackle the catastrophic overfitting problem, we introduce a Style Distillation Loss (SDL) module, which distills the noise images generated from all images with the noise images from individual image to decouple the style and content of images. The SDL module enables the diffusion model to concentrate on acquiring the stylistic attributes of images while reducing the influence of specific content of images. To mitigate catastrophic forgetting, we propose a Dual Rregularization for shared-LoRA (DR-LoRA) module, which is designed to facilitate the smooth transfer of knowledge from old styles. A LoRA weight regularization and a noise graph regularization are devised to maintain the knowledge learned from old styles, which fully consider both aspects of weights and features. All the customization tasks share a single set of LoRA parameters with DR-LoRA module to minimize memory consumption. Additionally, to address the limitation posed by the limited parameters of LoRA, a Task-wise Token Learning (TTL) module is develops to learn a distinct token for each style, which enables generation of different styles. The tokens corresponding to past styles are stored to further mitigate the catastrophic forgetting problem. Finally, we demonstrate the validation of our MuseumMaker through extensive experiments, and conduct ablation studies to emphasize the contribution of each module in our MuseumMaker.
@@ II-B Text to Image Generation
Other advanced works explore diffusion models in latent space, rather than operating directly on pixel space. Stable Diffusion[1] is trained with a large amount of data based on latent diffusion model (LDM)[34], which employs a VQ-GAN[35] for latent representation and adds text as conditional information to the denoising process. DALL-E 2[36] inverts the CLIP[37] image encoder with a diffusion model, learning a prior to connect the text and the features of images in the latent space. This text-image latent prior is found to be crucial for performance.
Building upon these pioneering efforts, subsequent studies have sought to enhance T2I diffusion models in various ways. These improvements encompass model architectures[38, 39, 40], spatial control via sketches or spatio-textual representation[41][42], textual inversion for controlling novel concepts[43], and retrieval mechanisms[44, 45, 46] to handle out-of-distribution scenarios. To further tackle the issue of continual learning for diffusion, [47][48] achieve the continuous generation of personalized concepts. However, the methods for continuous style learning and image generation are still lack of research.
@@ III-A Revisiting Text-to-Image Diffusion Model
The diffusion model [34] represents a probabilistic approach utilized for image generation, which generates a image by gradually denoising a noise graph from a Gaussian distribution. As exemplified by the Stable Diffusion (SD) technique [1], the primary objective of the diffusion model is to align semantic information between natural language and images throughout the training phase. Specifically, SD relies on a pre-trained text encoder ${\psi}$ from CLIP[37], a latent encoder $\mathcal{F}$ and a denoising U-Net $\epsilon_{\theta}$. Given a text prompt $p$, a timestep number $T$ and a random Gaussian noise graph $\varepsilon\in\mathcal{N}(0,\mathbf{I})$, the text embedding can be computed as $c=\psi(p)$. Formally, the training process of SD can be defined as follow:
@@
where timestep $t\in\text{Uniform}(1,T)$, $z_{t}$ is the image latents encoded by $\mathcal{E}$ at timestep $t$, and $\theta$ represents the paremeters of denosing model.
@@ III-B Problem Definition
Suppose that a series of continuous tasks for Text-to-Image (T2I) diffusion model are denoted as $\mathbf{D}^{K}=\{\mathcal{D}^{1}_{n_{k}},\mathcal{D}^{2}_{n_{k}},\cdots,\mathcal{D}^{K}_{n_{k}}\}$, where $K$ represents the number of consecutive tasks and $n_{k}$ denotes the number of samples per task. Consequently, we can define all the tasks requiring generation as $\mathcal{T}=\{\mathcal{T}^{k}\}_{k=1}^{K}$. For the $k$-th learning task, $\mathcal{T}^{k}$ corresponds to a dataset $\mathcal{D}^{k}_{n_{k}}=\{x^{k}_{i},p_{i}^{k}\}^{n_{k}}_{i=1}$, where $x^{k}_{i}\in\mathcal{X}^{k}$ and $p_{i}^{k}\in\mathcal{P}^{k}$ denote the $i$-th image and its corresponding prompt. Different from existing works[56][57], we consider a scenario that users provide consecutive images of different specific styles for T2I diffusion model, namely task $\mathcal{T}^{k}$. For each task, the number of pairs of image $x^{k}_{i}$ and prompt $p^{k}_{i}$ with $n_{k}$ are imposed into the continual style customization diffusion model, MuseumMaker. Considering the problem of privacy and memory limitation, the dataset of past tasks is unavailable. In this scenario, the model we proposed is able to incorporate the new style while ensuring the memory of past encountered styles. In details, the problem of continual style customization can be formulated as follows:
@@ IV-C DR-LoRA
$\mathcal{L}_{w}&=\frac{1}{L}\sum_{l=1}^{L}\left[1-\text{Sim}\left(\text{Flatten}(\Delta{\mathbf{W}_{l}^{k-1}}),\text{Flatten}(\Delta{\mathbf{W}_{l}^{k}})\right)\right],\end{aligned}$
@@ IV-D Task-wise Token Learning
However, Textual Inversion [63] overlooks the independence of cross-attention layers at different resolutions during denoising process. Solely training the input token may not sufficiently unfold the embedding space. Inspired by this observation, we extend the input token embedding by introducing multiple independent tokens, which are learned across different cross-attention layers of the U-Net. We denote the extended $v_{*}^{k}$ as $\mathcal{V}_{*}^{k}=\{v^{k}_{1},v^{k}_{2},...,v^{k}_{L}\}$, where $k$ represents the corresponding task index, $L$ represents the number of layers of cross-attention. The trainable tokens set of $K$ tasks can be written as $\boldsymbol{V}_{*}^{K}=\{\mathcal{V}_{*}^{1},\mathcal{V}_{*}^{2},...\mathcal{V}_{*}^{K}\}$. Consequently, the Eq. (9) for $K$ tasks can be rewritten as:
@@
Figure 4: Qualitative comparison of our ablation studies, where we evaluate the contribution of each module we proposed. We denotes the task-wise token learning module as TTL and style distillation loss as SDL in the fig, respectively. We input the same text prompt as we do in comparison experiments.
@@ V-A Datasets and Evaluation
V-B Implementation Details
In this work, we conduct continual style customization experiments on our proposed MuseumMaker, LoRA[61] and Specialist Diffusion (SPD)[56]. For the continual style customization experiment, we conduct the experiment on the 10 style datasets, ensuring that the previous style is not available during the training process. To assess the effectiveness of our continual style customization method in contrast to existing approaches, we deploy Fine-tuning (FT), as well as two classic continual learning methods, namely EWC [10] and LWF [68], for LoRA and SPD. These methods serve as benchmarks for comparison with our approach. Furthermore, we store a learnable token embedding and LoRA weight for each style as our upper bound. During the learning phase for each style, we employ 1000 training steps for LoRA baselines method, 100 training epoch for SPD baselines methods, 1000 training steps for ours. All the methods are implied with a batch size of 1. To achieve desirable generation performance, we set the learning rate to $1\times 10^{-5}$ for LoRA baselines method, $1\times 10^{-6}$ for SPD baselines methods and $1\times 10^{-5}$ for our method. Additionally, the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are set 0.8 and 1.0 in Eq. (8), $\alpha$ and $\beta$ in Eq. (12) are set to 0.8 and 1.5. During the inference stage, we apply a DDPM sampler with 50 steps for all methods. All experiments are conducted on a single NVIDIA RTX A6000 GPU, with all generated images having a resolution of $512\times 512$ pixels and based on Stable Diffusion v1.5.

### v2

@@ Abstract
Pre-trained large text-to-image (T2I) models with an appropriate text prompt has attracted growing interests in customized images generation field. However, catastrophic forgetting issue make it hard to continually synthesize new user-provided styles while retaining the satisfying results amongst learned styles. In this paper, we propose MuseumMaker, a method that enables the synthesis of images by following a set of customized styles in a never-end manner, and gradually accumulate these creative artistic works as a Museum. When facing with a new customization style, we develop a style distillation loss module to extract and learn the styles of the training data for new image generation. It can minimize the learning biases caused by content of new training images, and address the catastrophic overfitting issue induced by few-shot images. To deal with catastrophic forgetting amongst past learned styles, we devise a dual regularization for shared-LoRA module to optimize the direction of model update, which could regularize the diffusion model from both weight and feature aspects, respectively. Meanwhile, to further preserve historical knowledge from past styles and address the limited representability of LoRA, we consider a task-wise token learning module where a unique token embedding is learned to denote a new style. As any new user-provided style come, our MuseumMaker can capture the nuances of the new styles while maintaining the details of learned styles. Experimental results on diverse style datasets validate the effectiveness of our proposed MuseumMaker method, showcasing its robustness and versatility across various scenarios.
@@ I Introduction
Although most recent models[8][9] focus on boosting their generative capabilities with natural languages, these models fail to control the exact texture and color palettes of generation images, when adding popular or customized styles (e.g., “impressionist”) to the input text prompt. A naive way is to fine-tune the T2I diffusion model using few images of the given “impressionist” style, whereas this manner cannot synthesize various images of different specific styles in a never-ending manner. For example, as shown in Fig. 1, when the user inputs a prompt such as “a cat wearing sunglasses in the style of impressionism” to generate an impressionist image of the cat, a naive way is to fine-tune the diffusion model with images in impressionist style. Subsequently, if the user tries to augment this diffusion model with a new style, e.g., “cubism”, a simple method is to fine-tune the diffusion model with both impressionistic and cubist images. As users continuously seek to incorporate new styles, the large computational requirements and ever-increasing training times swiftly become impractical. Moreover, there is a high probability for diffusion model to overfit to the new user-provided styles after continual fine-tuning operator, resulting in unsatisfactory artistic generation performance[10].
@@
Inspired by the aforementioned considerations, we present a continual style learning approach based on T2I diffusion model in this work. Our proposed method aims to enable the continuous learning of new styles for generation purposes without accessing past datasets, which is designed to reduce memory consumption and compute efficiently. To address the two aforementioned challenges, we develop a continual style customization for diffusion model i.e, MuseumMaker, which could continuously adapt upcoming new styles and purely learn stylistic features from user-provided data. To tackle the catastrophic overfitting problem, we introduce a Style Distillation Loss (SDL) module, which decouples the style and content of images by extracting the representation of styles from the whole dataset. The SDL module enables the diffusion model to concentrate on acquiring the stylistic attributes of images while reducing the influence of specific content of images. To mitigate catastrophic forgetting, we propose a Dual Rregularization for shared-LoRA (DR-LoRA) module, which is designed to facilitate the smooth transfer of knowledge from old styles. This module incorporates LoRA weight regularization and stylistic feature regularization to preserve knowledge acquired from previous styles. All the customization tasks share a single set of LoRA parameters with DR-LoRA module to minimize memory consumption. Additionally, to address the limitation posed by the limited parameters of LoRA, a Task-wise Token Learning (TTL) module is develops to learn a distinct token for each style, which enables generation of different styles. The tokens corresponding to past styles are stored to further mitigate the catastrophic forgetting problem. Finally, we demonstrate the validation of our MuseumMaker through extensive experiments, and conduct ablation studies to emphasize the contribution of each module in our MuseumMaker.
@@ II-B Text to Image Generation
Other advanced works explore diffusion models in latent space, rather than operating directly on pixel space. Stable Diffusion[1] is trained with a large amount of data based on latent diffusion model (LDM), which employs a VQ-GAN[35] for latent representation and adds text as conditional information to the denoising process. DALL-E 2[36] inverts the CLIP[37] image encoder with a diffusion model, learning a prior to connect the text and the features of images in the latent space. This text-image latent prior is found to be crucial for performance.
Building upon these pioneering efforts, subsequent studies have sought to enhance T2I diffusion models in various ways. These improvements encompass model architectures[38, 39, 40], spatial control via sketches or spatio-textual representation[41][42][43], textual inversion for controlling novel concepts[44], and retrieval mechanisms[45, 46, 47] to handle out-of-distribution scenarios. To further tackle the issue of continual learning for diffusion, [48][49] achieve the continuous generation of personalized concepts. However, the methods for continuous style learning and image generation are still lack of research.
@@ III-A Revisiting Text-to-Image Diffusion Model
The diffusion model [58] represents a probabilistic approach utilized for image generation, which generates a image by gradually denoising a noise graph from a Gaussian distribution. As exemplified by the Stable Diffusion (SD) technique [1], the primary objective of the diffusion model is to learn a trajectory to generate the well patterned and distributed data from a random noise graph. Specifically, SD relies on a pre-trained text encoder ${\psi}$ from CLIP[37], a latent encoder $\mathcal{F}$, a decoder $\mathcal{G}$ and a denoising U-Net $\epsilon_{\theta}$. Given a text prompt $p$, a timestep $t\in\text{Uniform}(1,T)$ and a random Gaussian noise graph $\varepsilon\in\mathcal{N}(0,\mathbf{I})$, the text embedding can be computed as $c=\psi(p)$ ,and subsequently, the image $\hat{x}$ can be generated by $\hat{x}=\mathcal{G}(\epsilon_{\theta}(c,t))$. Formally, the training process of SD can be defined as follow:
@@
where $z_{t}$ is the image latents encoded by $\mathcal{E}$ at timestep $t$, and $\theta$ represents the parameters of denoising model.
@@ III-B Problem Definition
Suppose that a series of continuous tasks for Text-to-Image (T2I) diffusion model are denoted as $\mathbf{D}^{K}=\{\mathcal{D}^{1}_{n_{1}},\mathcal{D}^{2}_{n_{2}},\cdots,\mathcal{D}^{K}_{n_{K}}\}$, where $K$ represents the number of consecutive tasks and $n_{k}$ denotes the number of samples per task. Consequently, we can define all the generation tasks as $\mathcal{T}=\{\mathcal{T}^{k}\}_{k=1}^{K}$. For the $k$-th learning task, $\mathcal{T}^{k}$ corresponds to a dataset $\mathcal{D}^{k}_{n_{k}}=\{x^{k}_{i},p_{i}^{k}\}^{n_{k}}_{i=1}$, where $x^{k}_{i}\in\mathcal{X}^{k}$ and $p_{i}^{k}\in\mathcal{P}^{k}$ denote the $i$-th image and its corresponding prompt. Different from existing works[57][59], we consider a scenario that users provide consecutive images of different specific styles for T2I diffusion model, namely task $\mathcal{T}^{k}$. For each task, the number of pairs of image $x^{k}_{i}$ and prompt $p^{k}_{i}$ with $n_{k}$ are imposed into the continual style customization diffusion model, MuseumMaker. Considering the problem of privacy and memory limitation, the dataset of past tasks is unavailable. In this scenario, the model we proposed is able to incorporate the new style while ensuring the memory of past encountered styles. In details, the problem of continual style customization can be formulated as follows:
@@ IV-C DR-LoRA
$\mathcal{L}_{w}&=\frac{1}{L}\sum_{l=1}^{L}\left[1-\mathrm{Sim}\left(\text{Flatten}(\Delta{\mathbf{W}_{l}^{k-1}}),\text{Flatten}(\Delta{\mathbf{W}_{l}^{k}})\right)\right],\end{aligned}$
@@ IV-D Task-wise Token Learning
However, Textual Inversion [65] overlooks the independence of cross-attention layers at different resolutions during denoising process. Solely training the input token may not sufficiently unfold the embedding space. Inspired by this observation, we extend the input token embedding by introducing multiple independent tokens, which are learned across different cross-attention layers of the U-Net. We denote the extended $v_{*}^{k}$ as $\mathcal{V}_{*}^{k}=\{v^{k}_{1},v^{k}_{2},...,v^{k}_{L}\}$, where $k$ represents the corresponding task index, $L$ represents the number of layers of cross-attention. The trainable tokens set of $K$ tasks can be written as $\boldsymbol{V}_{*}^{K}=\{\mathcal{V}_{*}^{1},\mathcal{V}_{*}^{2},...,\mathcal{V}_{*}^{K}\}$. Consequently, the Eq. (9) for $K$ tasks can be rewritten as:
@@
Figure 4: Qualitative comparison of our ablation studies, where we evaluate the contribution of each module we proposed. We denotes the task-wise token learning module as TTL and style distillation loss as SDL in the fig, respectively. We input the same text prompt as we do in comparison experiments. The upper bound setting stores a learnable token embedding and LoRA weight for each style
@@ V-A Datasets and Evaluation
V-B Implementation Details and Baselines
In this work, we conduct continual style customization experiments on our proposed MuseumMaker, LoRA[63] and Specialist Diffusion (SPD)[57]. For the continual style customization experiment, we conduct the experiment on the 10 style datasets, ensuring that the previous style is not available during the training process. To assess the effectiveness of our continual style customization method in contrast to existing approaches, we deploy Fine-tuning (FT), as well as two classic continual learning methods, namely EWC [11] and LWF [70], for LoRA and SPD. These methods serve as benchmarks for comparison with our approach. Furthermore, we store a learnable token embedding and LoRA weight for each style as our upper bound. During the learning phase for each style, we employ 1000 training steps for LoRA baselines method, 100 training epoch for SPD baselines methods, 1000 training steps for ours. All the methods are implied with a batch size of 1. To achieve desirable generation performance, we set the learning rate to $1\times 10^{-5}$ for LoRA baselines method, $1\times 10^{-6}$ for SPD baselines methods and $1\times 10^{-5}$ for our method. Additionally, the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are set 0.8 and 1.0 in Eq. (8), $\alpha$ and $\beta$ in Eq. (12) are set to 0.8 and 1.5. During the inference stage, we apply a DDIM sampler with 50 steps for all methods. All experiments are conducted on a single NVIDIA RTX A6000 GPU, with all generated images having a resolution of $512\times 512$ pixels and based on Stable Diffusion v1.5.